Find the equation of the line which passes through the point (9,5) and has a slope 1

Give your answer in the form y=m x+b

The constants should be given exactly, e.g., 1/3 rather than 0.33.


1
Expert's answer
2021-10-24T01:28:20-0400

Let y=kx+by=kx+b - equation of the given line


k is a slope, and it is equal to 1


y=x+by=x+b


To find the value of, we can substitute x = 9.5, y = 9.5 since we know the line passes through this point


9.5=9.5+b \to b = 09.5=9.5+b→b=0